# Adaptive action supervision in reinforcement learning from real-world multi-agent demonstrations

### 15 Dec 2023 v3

@@ 1 INTRODUCTION
Modeling real-world biological multi-agents is a fundamental problem in various scientific and engineering fields. For example, animals, vehicles, pedestrians, and athletes observe others’ states and execute their own actions in complex situations. Pioneering works have proposed rule-based modeling approaches such as in human pedestrians [Helbing95] and animal groups [Couzin02] for each domain using hand-crafted functions (e.g., social forces). Recent advances in reinforcement learning (RL) with neural network approaches have enabled flexible and diverse modeling of such behaviors often in cyberspace [ross2010efficient, ho2016generative, hester2018deep].
However, when modeling real-world biological multi-agents, domain gaps may occur between behaviors in the sources (real-world data) with unknown dynamics and targets (cyberspace in RL) as shown in Fig. 1. The opposite configuration of the source and target has been actively studied and known as Sim-to-Real [rusu2017sim], which transfers the knowledge from cyberspace or human demonstrations to almost known source dynamics (simulation in Sim-to-Real) such as real-world robotics [schaal1996learning, kolter2007hierarchical]. In contrast, domain adaptation in real-world situations where the parameters of the source environment are often unknown cannot utilize explicit dynamics regarding source environments (e.g. transition model). In other words, we consider a Real-to-Sim domain adaptation problem in which the unknown source and the target are real-world data in a physical space and simulated data in cyberspace, respectively (Fig. 1). Team sports (e.g., football) and biological multi-agent motions (e.g., chase-and-escape) are examples that can be addressed with the above approach. Agents observe others’ states and execute planned actions [Fujii15a, Fujii16], but particularly in team sports, most of the governing equations are unknown. In such complex real-world behaviors, studies have separately investigated RL in cyberspace for flexible adaptation to complex environments [kurach2020google, li2021celebrating, tsutsui2022emergence] and data-driven modeling for reproducing real-world behaviors [Zheng16, Le17, fujii2021learning]. However, given the gap between these forward and backward approaches (see Fig. 1) in multi-agent RL (MARL) scenarios, an integrated approach to combine both strengths will be required.
@@
In this paper, we propose a method for adaptive action supervision in RL from real-world multi-agent demonstrations. To utilize the information of unknown source dynamics, we adopt an approach that combines RL and supervised learning by selecting the action of demonstrations based on the minimum distance between trajectories in source and target environments. Our goal is to balance reproducibility as imitations and generalization ability to obtain rewards (e.g., when different initial values are given in the same environment). These are mutually independent in general, and our approach that combines RL and supervised learning will help us achieve our goal. Compared with the case of Sim-to-Real in robotics, there have been no simulators for real-world agents with explicit dynamics in biological multi-agent scenarios. Therefore, in our experiments, we use simple chase-and-escape and football tasks with different dynamics between unknown source and target environments.
@@
In summary, our main contributions are as follows. (1) We propose a novel method for adaptive action supervision in RL from multi-agent demonstration, which will bridge the gap between RL in cyberspace and real-world data-driven modeling. (2) We adopt an approach that combines RL and supervised learning by selecting actions of demonstrations in RL based on the minimum distance of dynamic time warping (DTW) [vintsyuk1968speech] to utilize the information of the unknown source dynamics. This approach can be easily applied to many existing neural network architectures and provide us with an RL model balanced between reproducibility as imitation and generalization ability. (3) In the experiments, using a chase-and-escape and football tasks with the different dynamics between the unknown source and target environments, we show that our approach struck a balance between the reproducibility and generalization compared with the baselines. In particular, we used the tracking data of professional football players as expert demonstrations in a football RL environment and show successful performances. Our framework can estimate values for real agent behaviors and decision making if the model can imitate behaviors of players, which may be difficult for either data-driven and RL approaches.
@@
In the remainder of this paper, we describe the background of our problem and our method in Sections 2 and 3, overview related works in Section 4, and present experimental results and conclusions in Sections 5 and 6.
@@ 2 BACKGROUND
In this section, we consider a sequential decision-making setting in which multiple agents interact with a fully observable environment. Here we consider a forward RL model in Fig. 1 right. It is defined as a tuple $(K,S,A,\mathcal{T},R,\gamma)$, where $K$ is the fixed number of agents; $S$ is the set of states $s$; $A=[A_{1},...,A_{K}]$ represents the set of joint action $\vec{a}\in A$ (for a variable number of agents), and $A_{k}$ is the set of local action $a_{k}$ that agent $k$ can take; $\mathcal{T}(s^{\prime}|s,\vec{a}):S\times A\times S\rightarrow[0,1]$ is the transition model for all agents; $R=[R_{1},...,R_{K}]:S\times A\rightarrow\mathbb{R}^{K}$ is the joint reward function; and $\gamma\in(0,1]$ is the discount factor. In on-policy RL, the agent learns a policy $\pi_{k}:S_{k}\times A\rightarrow[0,1]$, where $S_{k}$ is a set of states for agent $k$.
The objective of agent $k$ is to discover the policy $\pi_{k}$ that maps states to actions, thereby maximizing the expected total reward over the agent’s lifespan, i.e., $G_{k}=\sum_{t}^{T}\gamma^{t}R_{k,t}$, where $R_{k,t}$ is the reward of agent $k$ at time $t$ and $T$ is the time horizon. The value $Q^{\pi}_{k}(s_{k},a_{k})$ related to a specific state-action pair $(s_{k},a_{k})$ serves as a expected future reward that can be acquired from $(s_{k},a_{k})$ when adhering to policy $\pi_{k}$. The optimal value function $Q^{*}(s,a)$, which offers the maximal values across all states, is determined by solving the Bellman equation:
@@
where $\mathcal{T}_{k}$ is the transition model of agent $k$. The optimal policy $\pi_{k}$ is then $\pi_{k}(s_{k})=\operatorname*{arg\,max}_{a_{k}\in A}Q_{k}^{*}(s_{k},a_{k}).$ Since our approach can be easily applied to many existing neural network architectures in model-free RL, we consider both independent policy for each agent and the joint policy $\vec{\pi}$ inducing the joint action-value function $Q_{tot}^{\vec{\pi}}(s,\vec{a})=\mathbb{E}_{s_{0:\infty},\vec{a}_{0:\infty}}\left[\sum_{t=0}^{\infty}\gamma^{t}R_{t}\mid s_{0}=s,\vec{a}_{0}=\vec{a},\vec{\pi}\right]$, where $R_{t}$ is the value of the joint reward at time $t$.
In a multi-agent system in complex real-world environments (e.g., team sports), (i) transition functions are difficult to design explicitly. Instead, (ii) if we can utilize the demonstrations of expert behaviors (e.g., trajectories of professional sports players), we can formulate and solve our problem as a machine learning problem (e.g., learning from demonstration or imitation learning). In other words, if the problem falls in the case that satisfies the two conditions, (i) and (ii), learning from demonstration is a better option than a pure RL approach by constructing the environment without demonstrations. As shown in Fig. 1 right, we perform forward RL with a temporal transition model $\mathcal{T}(s^{\prime}|s,\vec{a})$, but there should be a discrepancy between $\mathcal{T}$ and $\mathcal{T}^{E}(s^{\prime E}|s^{E},\vec{a}^{E})$ in Fig.1 left. Since we cannot access the information of $\mathcal{T}^{E}$, our solution is the supervised learning using the observed joint action $\vec{a}$ and adapted $\vec{a}^{E}$ in cyberspace for the learning of the Q function (see also Section 3). Next, we introduce DQN framework according to the previous work [hester2018deep]. For simplicity, we describe the following explanation using a single-agent RL framework and omit the agent index $k$.
@@
where $\theta^{\prime}$ refers to the parameters of the target network, and $a^{\max}_{t+1}=\operatorname*{arg\,max}_{a_{t}}Q(s_{t+1},a_{t};\theta)$. The upward bias typically associated with regular Q-learning updates is reduced by separating the value functions employed for these two variables. For more effcient learning, for example, to sample more significant transitions more frequently from its replay buffer, prioritized experience replay [schaul2016prioritized] have been used.
@@ 3 ADAPTIVE ACTION SUPERVISION IN RL
where $\mathbf{a}_{t}^{E}\in\{0,1\}^{|A|}$ (i.e., one-hot vector of expert actions), $\mathbf{q}_{s_{t}}=[Q(s_{t},a_{t}=1),...,Q(s_{t},a=|A|)]$, and the $\log$ applies element-wise. Ideally, Q-functions in the source and target domains should be compared, but when using limited real-world data, we consider that it would be better that more reliable action data is used as supervised data (rather than using approximated Q-function from data). This loss aims to achieve both reproducibility and generalization by maximizing the Q-function values for the action of the demonstration. A similar idea has been used [hester2018deep, Lakshminarayanan2016], which used only similar supervised losses in pre-training or RL, respectively, but we explicitly define and use this loss for both pre-training and RL to balance reproducibility and generalization.
Eq. (4) and the large margin classification loss in DQfD  [piot2014boosted] in Eq. (3) assumes that the timestamp of expert action $a^{E}_{t}$ should be the same as that of the RL model $a_{t}$. However, when there is a discrepancy between the source and target environments, the appropriate timestamp of expert actions can vary from that of the RL model actions. Thus, we propose a dynamic time-warping supervised loss for actions, which does not require prior knowledge, utilizing DTW framework [vintsyuk1968speech], which is a well-known algorithm in many domains [sakoe1978dynamic, myers1980performance, tappert1990state, kahveci2001variable, corradini2001dynamic, zhang2006comparison].
Here, we first consider two state sequences in RL and demonstration: $s=s_{1},\ldots,s_{t},\ldots,s_{n}$ and $s^{E}=s^{E}_{1},\ldots,s^{E}_{j},\ldots,s^{E}_{m}$, where $n$ and $m$ are the lengths of the sequences, respectively. We select $a^{E}_{t^{\prime}}$ at $t^{\prime}$ (which is not necessarily equal to $t$) for demonstration defined as:
@@ 4 RELATED WORK
In RL from demonstration [schaal1996learning], the direct approach recovers experts’ policies from demonstrations by supervised learning [pomerleau1991efficient, ross2010efficient, Ross11, sun2017deeply] or generative adversarial learning [ho2016generative, song2018multi], which make the learned policies similar to the expert policies (reviewed e.g., by [ramirez2022model] and [da2019survey, zhu2020transfer] as transfer learning). However, it is sometimes challenging to collect high-quality (e.g., optimal) demonstrations in many tasks. To obtain better policies from demonstrations, several methods combine imitation learning and RL [silver2016mastering, hu2018knowledge, hester2018deep, kang2018policy, Lakshminarayanan2016, nair2018overcoming, rajeswaran2017learning, zhang2018pretraining]. Some approaches [vecerik2017leveraging, hester2018deep] have been proposed to explore the sparse-reward environment by learning from demonstrations. Recently, cross domain adaptation problems have been considered to achieve the desired movements such as when changing morphologies [raychaudhuri2021cross, fickinger2022cross, franzmeyer2022learn]. In this case, imitation in terms of reproducibility would be difficult in principle because the problem (e.g., morphology) is changed. Our problem setting is different in terms of achieving both the ability to maximize a reward and reproducibility (imitation ability) rather than only the former.
@@
In RL applications, grid-world, robot Soccer, video games, and robotics have been intensively investigated. Among these domains, robotics and robot soccer are specifically related to real-world problems. In robotics, noise in sensors and actuators, limited computational resources, and the harmfulness of random exploration to people or animals are some of the many challenges [hua2021learning]. There have been successful applications of transfer learning in robotics [schaal1996learning, kolter2007hierarchical, sakato2014learning, isele2016using] (reviewed by e.g., [zhu2021survey]). These are mostly transferred from cyberspace or human demonstrations to real-world robotics (sometimes called Sim-to-Real [rusu2017sim]), which utilize almost known dynamics about the (at least) target dynamics. In contrast, our Real-to-Sim problem cannot utilize explicit dynamics about both source and target environments, and thus such domain adaptation in RL is challenging.
@@
Robot soccer is similar to our task, in particular, RoboCup (the Robot World Cup Initiative) involves attempts by robot teams to actually play a soccer game [Kitano97]. Some researchers have adopted imitation learning approaches [hussein2018deep, nguyen2020structure, wang2021multi], but the source and target environments are basically the same. In terms of simulators based on real-world data for data analysis, to our knowledge, there have been no domain adaptation methods in RL from real-world data to simulation environments.
@@
In the tactical behaviors of team sports, agents select an action that follows a policy (or strategy) in a state, receives a reward from the environment and others, and updates the state [fujii2021data]. Due to the difficulty in modeling the entire framework from data for various reasons [van2021learning] (e.g., a sparse reward and difficulty in estimating intents), we can adopt two approaches: to estimate the related variables and functions from data (i.e., inverse approach) as a sub-problem, and to build a model (e.g., RL model) to generate data in cyberspace (i.e., forward approach, e.g., [kurach2020google, li2021celebrating]).
@@
For the former, there have been many studies on inverse approaches. There have been many studies on estimating reward functions by inverse RL [Luo2020, Rahimian2020] and the state-action value function (Q-function) [Wang2018, Liu2018, Liu2020, ding2022deep]. Researchers have performed trajectory prediction in terms of the policy function estimation, as imitation learning [Le17, le2017data, Teranishi2020, fujii2020policy] and behavioral modeling [Zheng16, Zhan19, Yeh2019, Li2021, fujii2022estimating, teranishi2022evaluation] to mimic (not optimize) the policy using neural network approaches. This approach did not consider the reward in RL (and simulation) and usually performed a trajectory prediction.
@@
For the latter approach, researchers have proposed new MARL algorithms with efficient learning, computation, and communication [roy2020promoting, espeholt2019seed, liu2021semantic, liu2021towards, li2021celebrating, ruan2022gcs]. Recently, the ball-passing behaviors in artificial agents of Google Research Football (GFootball) [kurach2020google] and professional football players were compared [scott2022does], but a gap still exists between these forward and backward approaches. In other research fields, e.g., for animal behavioral analysis, forward [banino2018vector, ishiwaka2022deepfoids] and backward approaches [ashwood2022dynamic, fujii2021learning] have also been used separately. Our approach integrates both approaches to combine the reproducibility as imitation and generalization to obtain rewards.

### 19 Dec 2023 v4

@@ 1 INTRODUCTION
Modeling real-world biological multi-agents is a fundamental problem in various scientific and engineering fields. For example, animals, vehicles, pedestrians, and athletes observe others’ states and execute their own actions in complex situations. Pioneering works have proposed rule-based modeling approaches such as in human pedestrians [Helbing95] and animal groups [Couzin02] for each domain using hand-crafted functions (e.g., social forces). Recent advances in reinforcement learning (RL) with neural network approaches have enabled flexible and diverse modeling of such behaviors often in cyberspace [ross2010efficient, ho2016generative].
@@
However, when modeling real-world biological multi-agents, domain gaps may occur between behaviors in the sources (real-world data) with unknown dynamics and targets (cyberspace in RL) as shown in Fig. 1. The opposite configuration of the source and target has been actively studied and known as Sim-to-Real [rusu2017sim], which transfers the knowledge from cyberspace or human demonstrations to almost known source dynamics (simulation in Sim-to-Real) such as real-world robotics [schaal1996learning, kolter2007hierarchical]. In contrast, domain adaptation in real-world situations where the parameters of the source environment are often unknown cannot utilize explicit dynamics regarding source environments (e.g. transition model). In other words, we consider a Real-to-Sim domain adaptation problem in which the unknown source and the target are real-world data in a physical space and simulated data in cyberspace, respectively (Fig. 1). Team sports (e.g., football) and biological multi-agent motions (e.g., chase-and-escape) are examples that can be addressed with the above approach. Agents observe others’ states and execute planned actions [fujii2021data], and particularly in team sports, most of the governing equations are unknown. In such complex real-world behaviors, studies have separately investigated RL in cyberspace for flexible adaptation to complex environments [kurach2020google, li2021celebrating] and data-driven modeling for reproducing real-world behaviors [Zheng16, Le17, fujii2021learning]. However, given the gap between these forward and backward approaches (see Fig. 1) in multi-agent RL (MARL) scenarios, an integrated approach to combine both strengths will be required.
@@
In this paper, we propose a method for adaptive action supervision in RL from real-world multi-agent demonstrations. To utilize the information of unknown source dynamics, we adopt an approach that combines RL and supervised learning by selecting the action of demonstrations based on the minimum distance between trajectories in source and target environments. Our goal is to balance reproducibility as imitations and generalization ability to obtain rewards (e.g., when different initial values are given in the same environment). These are mutually independent in general, and our approach that combines RL and supervised learning will help us achieve our goal. Compared with the case of Sim-to-Real in robotics, there have been no simulators for real-world agents with explicit dynamics in biological multi-agent scenarios. In our experiments, we use simple chase-and-escape and football tasks with different dynamics between unknown source and target environments.
@@
In summary, our main contributions are as follows. (1) We propose a novel method for adaptive action supervision in RL from multi-agent demonstration, which will bridge the gap between RL in cyberspace and real-world data-driven modeling. (2) We adopt an approach that combines RL and supervised learning by selecting actions of demonstrations in RL based on the minimum distance of dynamic time warping (DTW) [vintsyuk1968speech] to utilize the information of the unknown source dynamics. This approach can be easily applied to existing neural network architectures and provide an RL model balanced between reproducibility as imitation and generalization ability. (3) In the experiments, using a chase-and-escape and football tasks with the different dynamics between the unknown source and target environments, our approach struck a balance between the reproducibility and generalization compared with the baselines. In particular, we used the tracking data of professional football players as expert demonstrations in a football RL environment and show successful performances. Our framework can estimate values for real agent behaviors and decision making if the model can imitate behaviors of players, which may be difficult for either data-driven and RL approaches.
@@
In the remainder, we describe the background of our problem and our method in Sections 2 and 3, overview related works in Section 4, and present results and conclusions in Sections 5 and 6.
@@ 2 BACKGROUND
Here, we consider a sequential decision-making setting of multiple agents interacting with a fully observable environment. We consider a forward RL model in Fig. 1 right, defined as a tuple $(K,S,A,\mathcal{T},R,\gamma)$, where $K$ is the fixed number of agents; $S$ is the set of states $s$; $A=[A_{1},...,A_{K}]$ represents the set of joint action $\vec{a}\in A$ (for a variable number of agents), and $A_{k}$ is the set of local action $a_{k}$ that agent $k$ can take; $\mathcal{T}(s^{\prime}|s,\vec{a}):S\times A\times S\rightarrow[0,1]$ is the transition model for all agents; $R=[R_{1},...,R_{K}]:S\times A\rightarrow\mathbb{R}^{K}$ is the joint reward function; and $\gamma\in(0,1]$ is the discount factor. In on-policy RL, the agent learns a policy $\pi_{k}:S_{k}\times A\rightarrow[0,1]$, where $S_{k}$ is a set of states for $k$.
The objective of agent $k$ is to discover the policy $\pi_{k}$ that maps states to actions, thereby maximizing the expected total reward over the agent’s lifespan, i.e., $G_{k}=\sum_{t}^{T}\gamma^{t}R_{k,t}$, where $R_{k,t}$ is the reward of agent $k$ at time $t$ and $T$ is the time horizon. The value $Q^{\pi}_{k}(s_{k},a_{k})$ related to a specific state-action pair $(s_{k},a_{k})$ serves as a expected future reward that can be acquired from $(s_{k},a_{k})$ when adhering to policy $\pi_{k}$. The optimal value function $Q^{*}(s,a)$, offering the maximal values across all states, is determined by the Bellman equation:
@@
where $\mathcal{T}_{k}$ is the transition model of agent $k$. The optimal policy $\pi_{k}$ is then $\pi_{k}(s_{k})=\operatorname*{arg\,max}_{a_{k}\in A}Q_{k}^{*}(s_{k},a_{k}).$ Since our approach can be easily applied to existing neural network in model-free RL, we consider both independent policy for each agent and the joint policy $\vec{\pi}$ inducing the joint action-value function $Q_{tot}^{\vec{\pi}}(s,\vec{a})=\mathbb{E}_{s_{0:\infty},\vec{a}_{0:\infty}}\left[\sum_{t=0}^{\infty}\gamma^{t}R_{t}\mid s_{0}=s,\vec{a}_{0}=\vec{a},\vec{\pi}\right]$, where $R_{t}$ is the value of the joint reward at time $t$.
In a multi-agent system in complex real-world environments (e.g., team sports), (i) transition functions are difficult to design explicitly. Instead, (ii) if we can utilize the demonstrations of expert behaviors (e.g., trajectories of professional sports players), we can formulate and solve it as a machine learning problem (e.g., learning from demonstration). In other words, if the problem falls in the case that satisfies the two conditions, (i) and (ii), learning from demonstration is a better option than a pure RL approach by constructing the environment without demonstrations. As shown in Fig. 1 right, we perform forward RL with a temporal transition model $\mathcal{T}(s^{\prime}|s,\vec{a})$, but there should be a discrepancy between $\mathcal{T}$ and $\mathcal{T}^{E}(s^{\prime E}|s^{E},\vec{a}^{E})$ in Fig.1 left. Since we cannot access the information of $\mathcal{T}^{E}$, our solution is the supervised learning using the observed joint action $\vec{a}$ and adapted $\vec{a}^{E}$ in cyberspace for the learning of the Q function (see also Section 3). Next, we introduce DQN framework according to the previous work [hester2018deep]. For simplicity, we describe the following explanation using a single-agent RL framework and omit the agent index $k$.
@@
where $\theta^{\prime}$ refers to the parameters of the target network, and $a^{\max}_{t+1}=\operatorname*{arg\,max}_{a_{t}}Q(s_{t+1},a_{t};\theta)$. The upward bias typically associated with regular Q-learning updates is reduced by separating the value functions employed for these two variables. For more efficient learning, e.g., to sample more significant transitions more frequently from its replay buffer, prioritized experience replay [schaul2016prioritized] have been used.
@@ 3 ADAPTIVE ACTION SUPERVISION IN RL
where $\mathbf{a}_{t}^{E}\in\{0,1\}^{|A|}$ (i.e., one-hot vector of expert actions), $\mathbf{q}_{s_{t}}=[Q(s_{t},a_{t}=1),...,Q(s_{t},a=|A|)]$, and the $\log$ applies element-wise. Ideally, Q-functions in the source and target domains should be compared, but when using limited data, it would be better that more reliable action data is used as supervised data (rather than using approximated Q-function from data). This loss aims to achieve both reproducibility and generalization by maximizing the Q-function values for the action of the demonstration. A similar idea has been used [hester2018deep, Lakshminarayanan2016], which used only similar supervised losses in pre-training or RL, respectively, but we explicitly define and use this loss for both pre-training and RL to balance reproducibility and generalization.
Eq. (4) and the large margin classification loss in DQfD  [piot2014boosted] in Eq. (3) assume that the timestamp of expert action $a^{E}_{t}$ should be the same as that of the RL model $a_{t}$. However, when there is a discrepancy between the source and target environments, the appropriate timestamp of expert actions can vary from that of the RL model actions. Thus, we propose a dynamic time-warping supervised loss for actions, which does not require prior knowledge, utilizing DTW framework [vintsyuk1968speech], a well-known algorithm in many domains [sakoe1978dynamic, myers1980performance, tappert1990state].
Here, we first consider two state sequences in RL and demonstration: $s=s_{1},\ldots,s_{t},\ldots,s_{n}$ and $s^{E}=s^{E}_{1},\ldots,s^{E}_{j},\ldots,s^{E}_{m}$, where $n$ and $m$ are the lengths of the sequences. We select $a^{E}_{t^{\prime}}$ at $t^{\prime}$ (which is not necessarily equal to $t$) for demonstration defined as:
@@ 4 RELATED WORK
In RL from demonstration [schaal1996learning], the direct approach recovers experts’ policies from demonstrations by supervised learning [pomerleau1991efficient, ross2010efficient, Ross11] or generative adversarial learning [ho2016generative, song2018multi], which make the learned policies similar to the expert policies (reviewed e.g., by [ramirez2022model] and [da2019survey, zhu2020transfer] as transfer learning). However, it is sometimes challenging to collect high-quality (e.g., optimal) demonstrations in many tasks. To obtain better policies from demonstrations, several methods combine imitation learning and RL such as [silver2016mastering, hu2018knowledge, Lakshminarayanan2016]. Some approaches [vecerik2017leveraging, hester2018deep] have been proposed to explore the sparse-reward environment by learning from demonstrations. Recently, cross domain adaptation problems have been considered to achieve the desired movements such as when changing morphologies [raychaudhuri2021cross, fickinger2022cross]. In this case, imitation in terms of reproducibility would be difficult in principle because the problem (e.g., morphology) is changed. Our problem setting is different in terms of achieving both the ability to maximize a reward and reproducibility (imitation ability) rather than only the former.
@@
In RL applications, grid-world, robot Soccer, video games, and robotics have been intensively investigated. Among these domains, robotics and robot soccer are specifically related to real-world problems. In robotics, noise in sensors and actuators, limited computational resources, and the harmfulness of random exploration to people or animals are some of the many challenges [hua2021learning]. There have been successful applications of transfer learning in robotics [schaal1996learning, kolter2007hierarchical, sakato2014learning] (reviewed by e.g., [zhu2021survey]).These are mostly transferred from cyberspace or human demonstrations to real-world robotics (sometimes called Sim-to-Real [rusu2017sim]), which utilize almost known dynamics about the (at least) target dynamics. In contrast, our Real-to-Sim problem cannot utilize explicit dynamics about both source and target environments, and thus such domain adaptation in RL is challenging.
@@
Robot soccer is similar to our task, in particular, RoboCup (the Robot World Cup Initiative) involves attempts by robot teams to actually play a soccer game [Kitano97]. Some researchers have adopted imitation learning approaches [hussein2018deep, nguyen2020structure], but the source and target environments are basically the same. In terms of simulators based on real-world data for data analysis, to our knowledge, there have been no domain adaptation methods in RL from real-world data to simulation environments.
@@
In the tactical behaviors of team sports, agents select an action that follows a policy (or strategy) in a state, receives a reward from the environment and others, and updates the state [fujii2021data]. Due to the difficulty in modeling the entire framework from data for various reasons [van2021learning], we can adopt two approaches: to estimate the related variables and functions from data (i.e., inverse approach) as a sub-problem, and to build a model (e.g., RL) to generate data in cyberspace (i.e., forward approach, e.g., [kurach2020google, li2021celebrating]).
@@
For the former, there have been many studies on inverse approaches. There have been many studies on estimating reward functions by inverse RL [Luo2020, Rahimian2020] and the state-action value function (Q-function) [Liu2018, Liu2020, ding2022deep, nakahara2023action]. Researchers have performed trajectory prediction in terms of the policy function estimation, as imitation learning [Le17, Teranishi2020, fujii2020policy] and behavioral modeling [Zheng16, Zhan19, Yeh2019, fujii2022estimating, teranishi2022evaluation] to mimic (not optimize) the policy using neural network approaches. This approach did not consider the reward in RL (and simulation) and usually performed a trajectory prediction.
@@
For the latter approach, researchers have proposed new MARL algorithms with efficient learning, computation, and communication [roy2020promoting, espeholt2019seed, liu2021semantic, li2021celebrating]. Recently, the ball-passing behaviors in artificial agents of Google Research Football (GFootball) [kurach2020google] and professional football players were compared [scott2022does], but a gap still exists between these forward and backward approaches. In other research fields, e.g., for animal behavioral analysis, forward [banino2018vector, ishiwaka2022deepfoids] and backward approaches [ashwood2022dynamic, fujii2021learning] have also been used separately. Our approach integrates both approaches to combine the reproducibility as imitation and generalization to obtain rewards.